# Cohomological Calibration and Curvature Constraints on Product Manifolds: A Topological Lower Bound

### v9

@@ Abstract.
We establish a quantitative relationship between mixed cohomology classes and the geometric complexity of metric connections with totally skew torsion on product manifolds. For any such connection whose non‑parallel torsion 33‑form represents a mixed de Rham class via its harmonic projection, we prove that the dimension of the off‑diagonal holonomy subspace satisfies the lower bound
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\;\geq\;\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K},$
where $\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$ is the direct sum of the kernels of the linear obstruction maps $\Psi_{p},\widetilde{\Psi}_{p}$. These kernels are defined intrinsically as the intersections
$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\quad\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr),$
where $\mathcal{V}_{2,1}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2})$ (respectively $\mathcal{V}_{1,2}\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2})$) denotes the linear span of the harmonic representatives of the $(2,1)$- (respectively $(1,2)$-) Künneth component of $\omega]$. The subspaces $\mathcal{V}_{2,1}$ and $\mathcal{V}_{1,2}$ are independent of the choice of a minimal harmonic decomposition. Consequently, the integer $\dim\mathcal{K}$ depends only on the mixed class $\omega]_{\mathrm{mixed}}$ and on the parallel‑form strata $\mathcal{P}_{k}(M_{i})$ of the factor metrics; hence the bound is invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these strata. This yields a computable, metric‑invariant obstruction to reducible holonomy and quantifies the topological “mixing strength” that forces irreducibility across the product splitting.
@@ 1. Introduction
More precisely, we prove
@@
where $\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$ is the direct sum of the kernels of the natural linear obstruction maps. To define these kernels, let
$\mathcal{V}_{2,1}:=\operatorname{span}\{\alpha_{i}\otimes\beta_{i}\}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2}),\qquad\mathcal{V}_{1,2}:=\operatorname{span}\{\widetilde{\alpha}_{j}\otimes\widetilde{\beta}_{j}\}\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2}),$
where $\{\alpha_{i}\otimes\beta_{i}\}$ (resp. $\{\widetilde{\alpha}_{j}\otimes\widetilde{\beta}_{j}\}$) is any minimal harmonic decomposition of the $(2,1)$- (resp. $(1,2)$-) Künneth component of $\omega]$. Because any two minimal harmonic decompositions span the same subspaces $U\subset\mathcal{H}^{2}(M_{1})$ and $V\subset\mathcal{H}^{1}(M_{2})$, the spaces $\mathcal{V}_{2,1}=U\otimes V$ and $\mathcal{V}_{1,2}$ are well‑defined, i.e., independent of the chosen decomposition. Then
The integer $\dim\mathcal{K}$ measures precisely the part of the mixed class that is absorbed by the parallel‑form strata and therefore does not contribute to off‑diagonal curvature. The quantity $r^{\sharp}$ is metric‑invariant under deformations that preserve the parallel‑form strata: it depends only on $\omega]_{\mathrm{mixed}}$ and on the subspaces $\mathcal{P}_{k}(M_{i})$. Consequently, each independent mixed Künneth tensor guarantees a linearly independent off‑diagonal holonomy direction, modulo the parallel forms captured by $\mathcal{K}$.
This yields a computable, purely topological criterion for forced irreducibility of the holonomy representation across the product splitting: irreducibility is forced whenever $r^{\sharp}>0$, i.e. whenever $\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})>\dim\mathcal{K}$.
@@ 3. Off-Diagonal Curvature and Holonomy Complexity
Definition 3.2 (Off-diagonal holonomy subspace).
Let $(M,g)$ be a Riemannian manifold and $\nabla^{C}$ a metric connection on $TM$ with totally skew–symmetric torsion. For a point $p\in M$, we define the off–diagonal holonomy algebra $\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})$ as the Lie subalgebra of $\mathfrak{hol}_{p}(\nabla^{C})$ generated by all off–diagonal components of the curvature tensor and its covariant derivatives at $p$, namely
$\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C}):=\mathrm{Lie}\,\Big\langle\big(\nabla^{k}R^{C}\big)_{p}(X_{1},\ldots,X_{k};Y,Z)^{\mathrm{off}}\;\Big|\;k\geq 0,\;X_{i},Y,Z\in T_{p}M\Big\rangle,$
where $(\cdot)^{\mathrm{off}}$ denotes the off–diagonal projection with respect to the product splitting $T_{p}M=V_{1}\oplus V_{2}$.
@@ 5. Main Result and Proof
Lemma 5.1 (Generic rank of covariant derivatives).
Let $(M_{1},g_{1})$ be a compact and oriented smooth Riemannian manifold and $\mathcal{H}^{2}(M_{1})$ the space of harmonic $2$-forms. Fix an $L^{2}$-orthonormal basis $\{\alpha_{i}\}_{i=1}^{r_{2,1}}$ of a subspace of $\mathcal{H}^{2}(M_{1})$ such that the first $s:=\dim\mathcal{P}_{2}$ elements span the parallel subspace $\mathcal{P}_{2}$. Then the set
$\mathcal{U}^{\prime}:=\Bigl\{p_{1}\in M_{1}\;\Big|\;\operatorname{rank}\!\bigl[(\nabla^{g_{1}}\alpha_{i})(p_{1})\bigr]_{i=s+1,\dots,r_{2,1}}=r_{2,1}-s\Bigr\}$
is open and dense in $M_{1}$.
@@
Fix smooth local frames of $\Lambda^{2}T^{*}M_{1}$ and $T^{*}M_{1}$ over an open coordinate ball $B\subset M_{1}$. In these frames each $(\nabla^{g_{1}}\alpha_{i})(p_{1})$ becomes a smooth $\mathbb{R}$-valued matrix $A_{i}(p_{1})$ of size $\dim M_{1}\times\binom{\dim M_{1}}{2}$. Form the $(r_{2,1}-s)\times\bigl[\dim M_{1}\cdot\binom{\dim M_{1}}{2}\bigr$] matrix
$\mathbf{A}(p_{1})=\bigl[A_{s+1}(p_{1})^{\top}\!,\dots,A_{r_{2,1}}(p_{1})^{\top}\bigr]^{\top}.$
The condition $\operatorname{rank}\mathbf{A}(p_{1})=r_{2,1}-s$ is equivalent to the non‑vanishing of at least one $(r_{2,1}-s)\times(r_{2,1}-s)$ minor of $\mathbf{A}(p_{1})$. Each minor is a smooth function on $B$; hence its non‑zero locus is open in $B$. Covering $M_{1}$ by such balls we conclude that $\mathcal{U}^{\prime}$ is open.
Now suppose, for a contradiction, that $\mathcal{U}^{\prime}$ is not dense. Then there exists a non‑empty open set $V\subset M_{1}$ on which
$\operatorname{rank}\mathbf{A}(p_{1})<r_{2,1}-s\qquad\forall\,p_{1}\in V.$
Since the rank is lower‑semicontinuous, after possibly shrinking $V$ we may assume that the rank is constant, say $\rho<r_{2,1}-s$, on $V$. Consequently, there exists a smooth map $\lambda:V\to\mathbb{R}^{r_{2,1}-s}\setminus\{0\}$ such that
$\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(p_{1})(\nabla^{g_{1}}\alpha_{i})(p_{1})=0\qquad\forall p_{1}\in V.$ (5.1)
Now we employ analytic approximation. By the theorem of Greene–Jacobowitz [4], we can approximate the metric $g_{1}$ in the $C^{\infty}$ topology by a sequence of real‑analytic metrics $\{g_{1}^{(k)}\}$. For each $k$, denote by $\alpha_{i}^{(k)}$ the $g_{1}^{(k)}$-harmonic representative of the cohomology class $\alpha_{i}]$. The forms $\alpha_{i}^{(k)}$ converge to $\alpha_{i}$ in $C^{\infty}$ as $k\to\infty$.
Because $\lambda$ is smooth and non‑zero on $V$, we may choose a point $p_{0}\in V$ and a subsequence (still denoted by $k$) such that $\lambda(p_{0})\neq 0$ and
$\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(p_{0})(\nabla^{g_{1}^{(k)}}\alpha_{i}^{(k)})(p_{0})=0\qquad\text{for all }k.$ (5.2)
Indeed, (5.1) passes to the limit after extracting a suitable subsequence, since all objects depend smoothly on the metric.
For each fixed $k$, the metric $g_{1}^{(k)}$ is real‑analytic. Define
$\eta^{(k)}:=\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(p_{0})\,\alpha_{i}^{(k)}\in\mathcal{H}^{2}(M_{1},g_{1}^{(k)}).$
Equation (5.2) states that $\nabla^{g_{1}^{(k)}}\eta^{(k)}(p_{0})=0$. Because $g_{1}^{(k)}$ is analytic, the coefficients of the Hodge–Laplacian are analytic, and Aronszajn’s unique-continuation theorem [1] applies. Since $\eta^{(k)}$ is harmonic and $g_{1}^{(k)}$ is real-analytic, $\eta^{(k)}$ is also real-analytic. The condition $\nabla_{g_{1}^{(k)}}\eta^{(k)}(p_{0})=0$ implies that $\eta^{(k)}$ is parallel in a neighborhood of $p_{0}$. By the unique-continuation principle for parallel forms on real-analytic manifolds (a consequence of Aronszajn’s theorem applied to the analytic coefficients), $\eta^{(k)}$ must be parallel on the entire connected component.
Passing to the limit $k\to\infty$, we obtain that
$\eta:=\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(p_{0})\,\alpha_{i}$
is parallel with respect to the original metric $g_{1}$ on the component containing $p_{0}$. Thus $\eta\in\mathcal{P}_{2}(M_{1})$, contradicting the linear independence of $\{\alpha_{s+1},\dots,\alpha_{r_{2,1}}\}$ modulo $\mathcal{P}_{2}(M_{1})$.
Therefore $\mathcal{U}^{\prime}$ must be dense in $M_{1}$.
The same reasoning applies verbatim to $\mathcal{H}^{k}(M_{i})$ for any $k$ and for both factors $M_{1}$ and $M_{2}$. ∎
@@
Let $M=M_{1}\times M_{2}$ be a compact and oriented product manifold with a product metric $g=g_{1}\oplus g_{2}$, and let $\nabla^{C}$ be a cohomologically calibrated metric connection with totally skew torsion $T$ such that $\omega]=[\omega_{h}]\in H^{3}(M;\mathbb{R})$ is a mixed class.
Let
$\mathcal{V}_{2,1}:=\operatorname{span}\{\alpha_{i}\otimes\beta_{i}\}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2}),\qquad\mathcal{V}_{1,2}:=\operatorname{span}\{\tilde{\alpha}_{j}\otimes\tilde{\beta}_{j}\}\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2}),$
where $\{\alpha_{i}\otimes\beta_{i}\}$ (resp. $\{\tilde{\alpha}_{j}\otimes\tilde{\beta}_{j}\}$) is any minimal harmonic decomposition of the $(2,1)$- (resp. $(1,2)$-) Künneth component of $\omega]$. Define
$\ker\Psi_{p}:=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\qquad\ker\widetilde{\Psi}_{p}:=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr),$
and set $\mathcal{K}:=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$. Then, for every $p\in M$,
@@
The bound depends only on the mixed class $\omega]_{\mathrm{mixed}}$ and on the parallel‑form strata $\mathcal{P}_{k}(M_{i})$; hence it is invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these strata.

### v10

@@ Abstract.
We establish a quantitative relationship between mixed de Rham classes and the geometric complexity of metric connections with totally skew torsion on admissible product manifolds: either both factors are compact oriented surfaces, or $\dim M=4$ with a $(3,1)$ splitting and the mixed class has tensor rank $1$. For any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion $T$, where the harmonic projection of $T$ represents a mixed class $\omega]$, we prove
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K},$
with $\mathcal{K}$ an intrinsically defined obstruction space. The bound is a topological invariant under metric deformations preserving the parallel‑form strata and provides a sharp obstruction to reducible holonomy. Counterexamples show the hypothesis is optimal.
@@ 1. Introduction
A subtle but crucial point is that the subspaces spanned by a minimal harmonic decomposition are uniquely determined only under topological restrictions on the factors. Consequently, our main result holds for admissible product manifolds (Definition 2.2), i.e.:
  • (A)
    both factors are compact oriented surfaces;  or
  • (B)
    $\dim M=4$ with a $(3,1)$ splitting and the mixed class has total tensor rank $1$.
For any admissible $M$, any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion satisfies
@@
where the obstruction space $\mathcal{K}$ is intrinsically defined via intersections of the mixed factor spaces $\mathcal{V}_{2,1}$, $\mathcal{V}_{1,2}$ with the parallel‑form strata $\mathcal{P}_{k}(M_{i})$. The bound is invariant under metric deformations preserving these strata.
When $r^{\sharp}>0$ the holonomy representation is forced to be irreducible across the product splitting. The examples in Section 6 illustrate both non‑trivial ($\Sigma_{g}\times T^{2}$) and trivial ($T^{2}\times T^{2}$) cases, and Remark 2.4 shows by explicit counterexample that without the admissibility hypothesis the bound fails to be well‑defined.
@@ 2. Cohomologically Calibrated Metric Connections with Skew Torsion
2.1. Intrinsic Definition of Mixed Cohomology Subspaces
We now address a subtle but crucial point: the subspaces spanned by the factors in a minimal harmonic decomposition are not uniquely determined in general. To obtain a well–defined geometric invariant we must impose conditions on the topology of the factors.
Definition 2.2 (Admissible Product Manifolds).
A compact oriented product manifold $M=M_{1}\times M_{2}$ is called admissible if it satisfies one of the following:
  • (A)
    both factors $M_{1},M_{2}$ are compact oriented surfaces;
  • (B)
    $\dim M=4$ with $(\dim M_{1},\dim M_{2})=(3,1)$ or $(1,3)$, and the mixed class $\omega]$ has tensor rank $\operatorname{rank}_{\mathbb{R}}([\omega]_{\text{mixed}})=1$.
Lemma 2.3 (Uniqueness of Mixed Subspaces).
Let $M$ be an admissible product manifold and $\omega]\in H^{3}(M;\mathbb{R})$ a non–trivial mixed class. Then the subspaces
$\mathcal{V}_{2,1}$ $\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2}),$ (2.7)
$\mathcal{V}_{1,2}$ $\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2})$ (2.8)
spanned by any minimal harmonic decomposition of the mixed Künneth components of $\omega]$ are intrinsically defined—i.e. they depend only on $\omega]$ and not on the particular minimal decomposition chosen.
Proof.
We treat the two admissible cases separately.
Case (A): $\dim M_{i}=2$. For a compact oriented surface $F^{2}$ we have $b_{2}(F)=1$ and $b_{1}(F)=2g(F)$. Let $\operatorname{vol}_{F}$ be the unique (up to scale) $g_{F}$–harmonic 22–form. Then
$\mathcal{H}^{2}(F)=\mathbb{R}\cdot\operatorname{vol}_{F}.$
Consequently any non–zero $T\in\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2})$ has rank one and can be written uniquely as
$T=\operatorname{vol}_{M_{1}}\otimes\beta,\qquad\beta\in\mathcal{H}^{1}(M_{2}).$
Indeed, if $\operatorname{vol}_{M_{1}}\otimes\beta=\operatorname{vol}_{M_{1}}\otimes\beta^{\prime}$, then $\operatorname{vol}_{M_{1}}\otimes(\beta-\beta^{\prime})=0$ and, because $\operatorname{vol}_{M_{1}}$ is non–zero as an element of $\mathcal{H}^{2}(M_{1})$, we must have $\beta=\beta^{\prime}$. Thus the one–dimensional subspace $\operatorname{span}\{\beta\}$ is uniquely determined by $T$, and
$\mathcal{V}_{2,1}=\mathcal{H}^{2}(M_{1})\otimes\operatorname{span}\{\beta\}$
is intrinsically defined. The same argument applied symmetrically yields the intrinsic definition of $\mathcal{V}_{1,2}$.
Case (B): $(3,1)$ or $(1,3)$ with $\operatorname{rank}_{\mathbb{R}}([\omega]_{\text{mixed}})=1$. Assume $(\dim M_{1},\dim M_{2})=(3,1)$ (the other case is symmetric). Then $b_{1}(M_{2})=1$ and $H^{2}(M_{2})=0$. Let $\gamma$ be a fixed basis of $\mathcal{H}^{1}(M_{2})$. Any $T\in\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2})$ has rank at most one and can be written uniquely as
$T=\alpha\otimes\gamma,\qquad\alpha\in\mathcal{H}^{2}(M_{1}).$
Uniqueness follows as in Case (A) because $\gamma$ is non–zero. Hence $\mathcal{V}_{2,1}=\operatorname{span}\{\alpha\}\otimes\mathcal{H}^{1}(M_{2})$ depends only on $\omega]$. The component $(1,2)$ is trivial since $H^{2}(M_{2})=0$, so $\mathcal{V}_{1,2}=\{0\}$ is obviously well defined. ∎
Remark 2.4 (Failure of invariance in general).
Invariance fails without the admissibility hypothesis. For instance, on $M=T^{3}\times T^{3}$ we have $b_{2}(T^{3})=3$ and $b_{1}(T^{3})=3$. The mixed class
$\omega]=(dx\wedge dy)\otimes dz+(dx\wedge dz)\otimes dy\in H^{2}(T^{3})\otimes H^{1}(T^{3})$
has rank $2$. It admits two distinct minimal decompositions
$\omega_{h}^{2,1}$ $=(dx\wedge dy)\otimes dz+(dx\wedge dz)\otimes dy$
$=\tfrac{1}{2}\big[(dx\wedge dy+dx\wedge dz)\otimes(dz+dy)$
$\qquad\quad+(dx\wedge dy-dx\wedge dz)\otimes(dz-dy)\big],$
which generate different subspaces $\mathcal{V}_{2,1}$ and $\mathcal{V}_{2,1}^{\prime}$. Consequently the integer $\dim\big(\mathcal{V}_{2,1}\cap(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2}))\big)$ depends on the choice, contradicting the metric–invariance.
2.2. Mixed obstruction spaces
Having established the intrinsic nature of the mixed factor spaces in Lemma 2.3 for admissible product manifolds (Definition 2.2), we now define the obstruction kernels.
Definition 2.5 (Mixed obstruction spaces).
For an admissible product manifold $M$ and a mixed class $\omega]$ we define
$\mathcal{V}_{2,1}$ $:=\operatorname{span}\{\alpha_{i}\otimes\beta_{i}\}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2}),$
$\mathcal{V}_{1,2}$ $:=\operatorname{span}\{\tilde{\alpha}_{j}\otimes\tilde{\beta}_{j}\}\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2}),$
where the spans are taken over any minimal harmonic decomposition of the mixed Künneth components of $\omega]$; by Lemma 2.3 these spans are independent of the particular choice.
The mixed obstruction spaces are the intersections
$\ker\Psi_{p}$ $:=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),$
$\ker\widetilde{\Psi}_{p}$ $:=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr),$
and we set $\mathcal{K}:=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$.
@@ 3. Off-Diagonal Curvature and Holonomy Complexity
Definition 3.2 (Off-diagonal holonomy algebra).
Let $(M,g)$ be a Riemannian product manifold with orthogonal splitting $T_{p}M=V_{1}\oplus V_{2}$, and let $\nabla^{C}$ be a metric connection on $TM$ with totally skew-symmetric torsion. Let $\Pi_{\mathrm{off}}:\mathfrak{so}(T_{p}M)\to\mathfrak{so}(T_{p}M)$ be the linear projection onto off-diagonal endomorphisms:
$\Pi_{\mathrm{off}}(A)(Z):=P_{1}A(P_{2}Z)+P_{2}A(P_{1}Z),$
where $P_{i}:T_{p}M\to V_{i}$ are the orthogonal projections.
Let $\mathcal{G}_{p}:=\{\nabla^{k}R^{C}|_{p}(X_{1},\dots,X_{k};Y,Z)\mid k\geq 0,X_{i},Y,Z\in T_{p}M\}$ be the set of Ambrose–Singer curvature generators of $\mathfrak{hol}_{p}(\nabla^{C})$. Define the off-diagonal curvature span
$\mathcal{S}_{p}^{\mathrm{off}}:=\operatorname{Span}_{\mathbb{R}}\{\Pi_{\mathrm{off}}(A)\mid A\in\mathcal{G}_{p}\}\subset\mathfrak{so}(T_{p}M).$
The off-diagonal holonomy algebra is the Lie subalgebra obtained by intersecting $\mathfrak{hol}_{p}(\nabla^{C})$ with the Lie algebra generated by $\mathcal{S}_{p}^{\mathrm{off}}$ in $\mathfrak{so}(T_{p}M)$:
$\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C}):=\bigl\langle\mathcal{S}_{p}^{\mathrm{off}}\bigr\rangle_{\mathrm{Lie}}\cap\mathfrak{hol}_{p}(\nabla^{C})\subset\mathfrak{hol}_{p}(\nabla^{C}).$
Here $\langle\cdot\rangle_{\mathrm{Lie}}$ denotes the Lie algebra generated in $\mathfrak{so}(T_{p}M)$, and the intersection ensures that $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$ is canonically a subalgebra of $\mathfrak{hol}_{p}(\nabla^{C})$. Equivalently, it is the smallest Lie subalgebra of $\mathfrak{hol}_{p}(\nabla^{C})$ whose off-diagonal components contain $\mathcal{S}_{p}^{\mathrm{off}}$.
@@ 5. Main Result and Proof
Lemma 5.1 (Existence of points of maximal rank for covariant derivatives).
Let $(M_{1},g_{1})$ be a compact oriented Riemannian manifold and let $\mathcal{H}^{2}(M_{1},g_{1})$ denote the space of harmonic $2$-forms. Given a subspace $U\subset\mathcal{H}^{2}(M_{1},g_{1})$ of dimension $r$ and its parallel part $U_{\parallel}:=U\cap\mathcal{P}_{2}(M_{1},g_{1})$ of dimension $s$, choose a basis $\alpha_{1},\dots,\alpha_{r}$ of $U$ such that $\alpha_{1},\dots,\alpha_{s}$ span $U_{\parallel}$. Then there exists a non‑empty open set $\mathcal{U}\subset M_{1}$ on which the covariant derivatives
$\bigl(\nabla^{g_{1}}\alpha_{s+1},\dots,\nabla^{g_{1}}\alpha_{r}\bigr)$
are pointwise linearly independent; i.e.
$\mathcal{U}=\Bigl\{p\in M_{1}\;\Big|\;\operatorname{rank}\bigl[(\nabla^{g_{1}}\alpha_{i})(p)\bigr]_{i=s+1,\dots,r}=r-s\Bigr\}\neq\emptyset.$
@@
We proceed in three steps.
1. Reduction to a real‑analytic metric. Real‑analytic Riemannian metrics are dense in the space of smooth metrics on a compact manifold; this follows from the theorem of Greene–Jacobowitz on analytic approximation of smooth maps (see [4]). Hence we can choose a sequence of real‑analytic metrics $\{g_{1}^{(k)}\}_{k\in\mathbb{N}}$ converging to $g_{1}$ in the $C^{\infty}$ topology. The dimension of the space of parallel $2$-forms is upper‑semicontinuous with respect to the metric (see [3]); therefore, after passing to a subsequence, we may assume
$\dim\mathcal{P}_{2}\bigl(M_{1},g_{1}^{(k)}\bigr)\leq\dim\mathcal{P}_{2}(M_{1},g_{1})=s\qquad\text{for all }k.$
For each $k$ let $\alpha_{i}^{(k)}\in\mathcal{H}^{2}\bigl(M_{1},g_{1}^{(k)}\bigr)$ be the unique harmonic representative of the cohomology class $\alpha_{i}]$. By the stability theorem for harmonic forms under metric deformation (Warner [9], see also Kodaira–Spencer [7]), $\alpha_{i}^{(k)}\to\alpha_{i}$ in $C^{\infty}$. Consequently the covariant derivatives converge in $C^{1}$:
$\nabla^{g_{1}^{(k)}}\alpha_{i}^{(k)}\longrightarrow\nabla^{g_{1}}\alpha_{i}\qquad(i=s+1,\dots,r).$
If we prove that for each $k$ there exists a non‑empty open set $\mathcal{U}^{(k)}\subset M_{1}$ where the $\nabla^{g_{1}^{(k)}}\alpha_{i}^{(k)}$ are linearly independent, then the $C^{1}$ convergence guarantees that, for large enough $k$, the same linear independence holds on a non‑empty open set for the original metric $g_{1}$. Indeed, linear independence is equivalent to the non‑vanishing of at least one $(r-s)\times(r-s)$ minor of the matrix of components; each minor depends continuously on the metric and the forms, so a minor that is non‑zero on an open set for $g_{1}^{(k)}$ remains non‑zero on an open set for $g_{1}$ when $k$ is sufficiently large.
Thus it suffices to prove the lemma under the assumption that $g_{1}$ is real‑analytic.
2. The real‑analytic case. Assume now that $g_{1}$ is real‑analytic. Then every harmonic form is real‑analytic (by elliptic regularity for the Hodge–Laplacian, see [8]), and consequently the covariant derivatives $\nabla^{g_{1}}\alpha_{i}$ are real‑analytic sections of the vector bundle $E:=T^{*}M_{1}\otimes\Lambda^{2}T^{*}M_{1}$.
Consider the real‑analytic section
$\Phi:M_{1}\longrightarrow\Lambda^{r-s}E,\qquad\Phi(p):=(\nabla^{g_{1}}\alpha_{s+1})(p)\wedge\cdots\wedge(\nabla^{g_{1}}\alpha_{r})(p).$
The set $\mathcal{U}:=\{p\in M_{1}\mid\Phi(p)\neq 0\}$ is exactly the set where the covariant derivatives are linearly independent; because $\Phi$ is continuous, $\mathcal{U}$ is open. We must show that $\mathcal{U}\neq\emptyset$.
Suppose, for contradiction, that $\mathcal{U}=\emptyset$; i.e. $\Phi(p)=0$ for every $p\in M_{1}$. Then at each point the sections $\nabla^{g_{1}}\alpha_{s+1},\dots,\nabla^{g_{1}}\alpha_{r}$ are linearly dependent. Since the condition of rank $\leq d$ is given by the vanishing of all $(d+1)\times(d+1)$ minors, which are analytic functions, the locus where the rank is constant is an analytic subset. Let d<rsd<r-s be the minimal rank and let $p_{0}$ be a point where the rank equals $d$. There exists a connected open neighbourhood $V$ of $p_{0}$ on which the rank is constant and equal to $d$.
On $V$ we can therefore find real‑analytic functions $c_{1},\dots,c_{r-s}:V\to\mathbb{R}$, not all zero, such that
$\sum_{i=1}^{r-s}c_{i}(p)\bigl(\nabla^{g_{1}}\alpha_{s+i}\bigr)(p)=0\qquad\forall p\in V.$ (5.1)
Define the analytic $2$-form $\eta:=\sum_{i=1}^{r-s}c_{i}\alpha_{s+i}$ on $V$. From (5.1) we have $\nabla^{g_{1}}\eta=\sum_{i=1}^{r-s}(dc_{i})\otimes\alpha_{s+i}+\sum_{i=1}^{r-s}c_{i}\nabla^{g_{1}}\alpha_{s+i}=\sum_{i=1}^{r-s}(dc_{i})\otimes\alpha_{s+i}$.
Let $V\subset M_{1}$ be the non–empty open subset fixed above. Since the mixed cohomology class on $M_{1}\times M_{2}$ is assumed to be non–trivial, its harmonic representative has a non–vanishing component in $H^{2}(M_{1})\otimes H^{1}(M_{2})$. In particular, the projection onto $H^{2}(M_{1})$ is non–zero. Hence the harmonic forms $\alpha_{s+1},\dots,\alpha_{r}$ cannot all vanish identically on $M_{1}$. Let $W\subset M_{1}$ denote the complement of their common zero set. By real–analyticity, $W$ is a non–empty open subset of $M_{1}$.
Restricting to the non‑empty open set $V^{\prime}:=V\cap W$, we have on $V^{\prime}$
$\sum_{i=1}^{r-s}(dc_{i})\otimes\alpha_{s+i}=0.$
Because the $\alpha_{s+i}(p)$ are linearly independent at every $p\in V^{\prime}$, this forces $dc_{i}=0$ on $V^{\prime}$ for each $i$. Thus the functions $c_{i}$ are constant on each connected component of $V^{\prime}$. By analyticity, they extend to constants on the whole connected open set $V$ (if two analytic functions coincide on an open set, they coincide everywhere on the connected component). Consequently, the $c_{i}$ are constant on $V$. Thus there exist constants $c_{1},\dots,c_{r-s}\in\mathbb{R}$, not all zero, such that
$\sum_{i=1}^{r-s}c_{i}\nabla^{g_{1}}\alpha_{s+i}=0\qquad\text{on }V.$
Define the global harmonic $2$-form
$\zeta:=\sum_{i=1}^{r-s}\bar{c}_{i}\alpha_{s+i}\in\mathcal{H}^{2}(M_{1}).$
Then $\nabla^{g_{1}}\zeta=0$ on $V$; i.e., $\zeta$ is parallel on the open set $V$. Because both $\zeta$ and the condition $\nabla^{g_{1}}\zeta=0$ are real-analytic, the unique-continuation principle [1] implies that $\nabla^{g_{1}}\zeta=0$ on the whole connected component of $M_{1}$ containing $V$, hence on all of $M_{1}$. Therefore $\zeta$ is a globally parallel $2$-form. But this contradicts the hypothesis that $\alpha_{s+1},\dots,\alpha_{r}$ are linearly independent modulo $\mathcal{P}_{2}(M_{1},g_{1})$. Indeed, $\zeta$ is a non-trivial linear combination of these forms that belongs to $\mathcal{P}_{2}(M_{1},g_{1})$. Hence our assumption that $\mathcal{U}=\emptyset$ is impossible, and we conclude that $\mathcal{U}$ is a non-empty open subset of $M_{1}$.
3. Returning to the original smooth metric. As explained in Step 1, the existence of a non‑empty open set of linear independence for a sequence of real‑analytic approximations $g_{1}^{(k)}$ together with the $C^{1}$ convergence of the covariant derivatives guarantees the existence of a non‑empty open set with the same property for the original smooth metric $g_{1}$. This completes the proof. ∎
@@
Let $M=M_{1}\times M_{2}$ be an admissible compact oriented Riemannian product manifold (Definition 2.2) with product metric $g=g_{1}\oplus g_{2}$. Let $\nabla^{C}$ be a cohomologically calibrated metric connection with totally skew torsion $T$ such that $\omega]=[\omega_{h}]$ is a non‑trivial mixed class.
Define the mixed factor subspaces
$\mathcal{V}_{2,1}:=\operatorname{span}\{\alpha_{i}\otimes\beta_{i}\}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2}),$
$\mathcal{V}_{1,2}:=\operatorname{span}\{\tilde{\alpha}_{j}\otimes\tilde{\beta}_{j}\}\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2}),$
where the spans are taken over any minimal harmonic decomposition of the mixed Künneth components of $\omega]$; in the admissible setting these spans are independent of the particular choice (Lemma 2.3).
Define the obstruction kernels
$\ker\Psi_{p}:=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\quad\ker\widetilde{\Psi}_{p}:=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr),$
and set $\mathcal{K}:=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$.
Then, for every $p\in M$,
@@
The bound depends only on $\omega]_{\text{mixed}}$ and on the parallel-form strata $\mathcal{P}_{k}(M_{i})$; hence it is invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these strata.